[figure: 9.7 cm]

THE INFLUENCE OF MECHANICAL ALLOYING ON THE PHYSICAL PROPERTIES OF NANOSTOICHIOMETRIC U-Fe

COMPOUNDS

M. Timko, A. Szlaferek^{a}, A. Zentko, J. Kováč and V. Kavečanský Institute of Experimental Physics, SAS, Watsonova 47, 043 53 Košice, Slovak Republic

aInstitute of Molecular Physics, PAS, Smoluchowskiego 17, 60-179 Poznań, Poland

It has been investigated the influence of mechanical milling on physical properties of UFe_{x} intermetallic compounds. The samples were prepared by a crucible- less, semi levitation, induction melting and then milled for appropriate time in a conventional horizontal ball mill. X-ray diffraction, differential scanning calorimetry and measurements of magnetization by vibrating sample magnetometer were performed to study of physical properties after milling. The X-ray patterns and DSC have indicated that the mechanical alloying (MA) results in a formation of an amorphous iron reach U-Fe phase. The observed values of T_{C} increase abruptly during the initial 65 h MA and then more slowly up to 207 K when the milling period rises up to 230 h.

[figure: 13.4 cm]

Subject category :

3. Transition Metals, Alloys and Compounds

Presentation mode : poster

Corresponding author : M. Timko

Address for correspondence :

Institute of Experimental Physics, SAS, Watsonova 47, 043 53 Košice, Slovak Republic

Email address : timko@ns.saske.sk